OpenStudy (anonymous):

Solve, using the substitution method. 3x + 2y = 11 x – 2 = – 4y A.(4,-1/2) B.(12,-5/2) C.(0,-2) D.(0,-1/2) I got D is that right?

OpenStudy (compassionate):

Hi, my name is Stephen. How are you, Ariarose? I see you like cute little girls. I think they're adorable, too (: 3x + 2y = 11 x - 2 = -4y So, we want to use elimination. What, when multiplied, will cancel out y? This is a tough one... if we look at substitution, we can add 2 to both sides and isolate x. 3x + 2y = 11 x = -4y + 2 Now, we plug in x... 3(-4y + 2) + 2y = 11 Solve for y. Once you have y, plug it into either equation to solve for x.

OpenStudy (anonymous):

add 2 to second equation . it becomes \[x=2-4y\] now put it into first equation. \[3 (2-4y) + 2y = 11\] \[6-12y + 2y =11 \implies 6-10y=11 \implies y = -\frac{ 1 }{ 2 }\] now put it into any equation to get \[x = 4\] so solution is \[ A(4, -\frac{ 1 }{ 2 })\]

OpenStudy (compassionate):

Good job, crazysignh. I appreciate your work. (:

OpenStudy (anonymous):

ohh ok Thank you both so much! :D I wish we could give 2 medals out i would give both of you one :D

OpenStudy (the_fizicx99):

Technically you \(\ could \) use elimination to solve this xD